# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ Appendix E Proofs in Section 5
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right].$
@@
$\phi(x)=\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\right]=\sum_{\left(d_{1},...,d_{I}\right)}{l\left(d_{1},d_{2},...,d_{I};x\right)\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right).$
@@
$=$ $\sum_{\left(d_{1},...,d_{I}\right)}{\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right)+\sum_{\left(d_{1},...,d_{I}\right)}{l\left(d_{1},d_{2},...,d_{I};x\right)\nabla\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right)$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\nabla\log\Pi_{i=1}^{I}P\left(d_{i};\pi^{*}\left(x\right)\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\nabla\left(\sum_{i}{\log P\left(d_{i};\pi^{*}\left(x\right)\right)}\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\log\pi_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)}}\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right],$
where the first equality comes from the chain rules, the second equality is obtained by “log probability tricks” in REINFORCE [75], and the last equality takes advantege of (5). ∎
@@ Appendix G Analysis of M-SoBiRL
In this section, we focus on the convergence analysis of Algorithm 1, M-SoBiRL. To begin with, a short proof sketch is provided for guidance.
@@
The proof is structured in four main steps.
Step1: preliminary properties.
@@
Step2: upper-bounding the residual $\left\|{w_{k}-w_{k}^{*}}\right\|_{2}$.
@@
$\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}\leq$ $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
where the coefficient $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}\right)$ implies its descent property, and the constants $C_{\sigma\pi},L_{w}$ will be specified later.
Step3: upper-bounding the errors $\left\|{Q_{k}-Q_{k}^{*}}\right\|_{\infty}$.
@@
Step4: Assembling the estimations above and achieving the conclusion.
@@ G.2 Lipschitz properties of quantities related to Algorithm 1
Under Assumptions 6.1 and 6.3, $\vartheta\left(x\right)$ is $L_{\vartheta}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@
Under Assumptions 6.1 and 6.3, $\nabla\phi(x)$ is $L^{M}_{\phi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@ G.3 Convergence property of $w_{k}$
Under Assumptions 6.1 and 6.3, it holds that
@@
Under Assumptions 6.1 and 6.3, the following inequalities hold in Algorithm 1.
@@
Under Assumption 6.1 and 6.3, the iterates $\{(x_{k},\pi_{k},w_{k})\}$ generated by Algorithm 1 satisfy
$\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}\leq$ $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
@@
By the update rule of $w$, $w_{k}=w_{k-1}-\xi\sigma_{k}$, we decompose the first term in (32):
$\left\|w_{k}-w_{k-1}^{*}\right\|_{2}^{2}=\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}-2\xi\left<\sigma_{k},w_{k-1}-w_{k-1}^{*}\right>.$ (33)
@@
$\xi C_{\sigma\pi}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}\leq$ $\frac{\xi^{2}}{2\rho_{1}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}+\frac{C_{\sigma\pi}^{2}\rho_{1}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}$
$=$ $\frac{\xi^{2}C_{\sigma\pi}^{2}}{2\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}+\frac{\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
@@
$\xi\left<\sigma_{k},w_{k-1}-w_{k-1}^{*}\right>\geq\frac{\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}-\frac{\xi^{2}C_{\sigma\pi}^{2}}{2\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}.$ (35)
@@
$\left\|w_{k}-w_{k-1}^{*}\right\|_{2}^{2}\leq\left(1-\xi\left(1-\gamma\right)^{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\frac{\xi^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}.$ (36)
@@
$\leq$ $\ \rho_{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{1}{\rho_{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}$
$=$ $\ \frac{\xi\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{2}{\xi\left(1-\gamma\right)^{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}$
$+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2},$ (38)
where we take $\,\,\rho_{2}=\frac{\xi\left(1-\gamma\right)^{2}}{2}$. Assembling (36), (37) and (38), we can estimate $\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}$ in (32),
@@
$\leq$ $\left(1-\xi\left(1-\gamma\right)^{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\frac{\xi^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+\left\|w_{k-1}^{*}-w_{k}^{*}\right\|_{2}^{2}+\frac{\xi\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}$
$+\frac{2}{\xi\left(1-\gamma\right)^{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}\,\,$
$=$ $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{\xi^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+2\xi^{2}\left\|\sigma_{k}\right\|_{2}^{2}$
$\leq$ $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\xi^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$

### v2

@@ Appendix C PRELIMINARIES ON NUMERICAL LINEAR ALGEBRA
We first revisit some basics for non-negative matrices and Markov chains; see (Seneta,, 2006).
@@ Appendix E PROOFS IN SECTION 5
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right].$
@@
$\phi(x)=\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]=\sum_{\left(d_{1},\ldots,d_{I}\right)}{l\left(d_{1},d_{2},\ldots,d_{I};x\right)\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right).$
@@
$=$ $\sum_{\left(d_{1},\ldots,d_{I}\right)}{\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right)+\sum_{\left(d_{1},\ldots,d_{I}\right)}{l\left(d_{1},d_{2},\ldots,d_{I};x\right)\nabla\Pi_{i=1}^{I}}P\left(d_{i};\pi^{*}\left(x\right)\right)$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\nabla\log\Pi_{i=1}^{I}P\left(d_{i};\pi^{*}\left(x\right)\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\nabla\left(\sum_{i}{\log P\left(d_{i};\pi^{*}\left(x\right)\right)}\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]+\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\log\pi_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)}}\right)\right]$
$=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right],$
where the first equality comes from the chain rules, the second equality is obtained by “log probability tricks” in REINFORCE (Sutton and Barto,, 2018), and the last equality takes advantege of (5). ∎
@@ Appendix G ANALYSIS OF M-SOBIRL
0:  outer and inner iteration numbers $K,N$, step sizes $\beta,\eta$, initialization $x_{0},\pi_{0},Q_{0},w_{0}$
1:  for $k=1,\dots,K$ do
2:     $A_{k}=\left({I-\gamma P^{\pi_{k}}}\right)^{\top}$$b_{k}=U^{\top}\operatorname{diag}\left({\pi_{k}}\right)\nabla_{\pi}f(x_{k},\pi_{k})$
3:     Compute $V_{k}$ by (9)
4:     $w_{k}=w_{k-1}-\eta\left({\left({A_{k}^{\top}A_{k}}\right)w_{k-1}-A_{k}^{\top}b_{k}}\right)$
6:     $Q_{k,0}=Q_{k}$
7:     for $n=1,\dots,N$ do
8:        $Q_{k,n}=\mathcal{T}_{\mathcal{M}_{\tau}(x_{k+1})}(Q_{k,n-1})$
9:     end for
11:  end for
11:  $(x_{K+1},\pi_{K})$
Enlightened by the hyper-gradient,
$-\tau^{-1}\nabla_{x}\varphi\left({x,V^{*}(x)}\right)^{\top}\left({I-\gamma P^{\pi^{*}(x)}}\right)^{-\top}U^{\top}\operatorname{diag}\left({\pi^{*}(x)}\right)\nabla_{\pi}f(x,\pi^{*}(x)),$ (24)
we propose the model-based algorithm, M-SoBiRL; see Algorithm 2. Note that in practice, the special structure of RL will facilitate the computation. Specifically, the matrix $A_{k}=(I-\gamma P^{\pi_{k}})^{\top}$ is usually sparse in RL, and the calculation of $b_{k}=U^{\top}\mathrm{diag}(\pi_{k})\nabla_{\pi}f(x_{k},\pi_{k})$ also involves a sparse constant matrix $U$ and a diagonal $\mathrm{diag}(\pi_{k})$. Therefore, the multiplications $A^{\top}_{k}A_{k}$ and $A_{k}^{\top}b_{k}$ for updating $w_{k}$ in line 4 can be accelerated by sparse multiplication oracles, the cost of which is denoted as $c_{\mathrm{sparse}}$. Additionally, evaluating ^$\hat{\nabla}\phi$ and updating $x$ in line 5 cost $\mathcal{O}(|\mathcal{S}||\mathcal{A}|n)$. Moreover, the other lines in Algorithm 2 totally cost $\mathcal{O}(|\mathcal{S}||\mathcal{A}|)$ since they are element-wise operations on $Q_{k}$ or $V_{k}$. In summary, the computation cost of each outer iteration is $\mathcal{O}(|\mathcal{S}||\mathcal{A}|n+c_{\mathrm{sparse}})$.
Next, we focus on the convergence analysis for M-SoBiRL, starting with a short proof sketch for guidance.
@@
The proof is structured into four main steps.
Step1: Preliminary Properties
@@
Step2: Upper-bounding the Residual $\left\|{w_{k}-w_{k}^{*}}\right\|_{2}$
@@
$\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}\leq$ $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
where the coefficient $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}\right)$ implies its descent property, and the constants $C_{\sigma\pi},L_{w}$ will be specified later.
Step3: Upper-bounding the Errors $\left\|{Q_{k}-Q_{k}^{*}}\right\|_{\infty}$
@@
Step4: Assembling the Estimations above and Achieving the Conclusion
@@ G.2 Lipschitz Properties of Quantities Related to Algorithm 2
Under Assumptions 7.1 and 7.3, $\vartheta\left(x\right)$ is $L_{\vartheta}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@
Under Assumptions 7.1 and 7.3, $\nabla\phi(x)$ is $L^{M}_{\phi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@ G.3 Convergence Property of $w_{k}$
Under Assumptions 7.1 and 7.3, it holds that
@@
Under Assumptions 7.1 and 7.3, the following inequalities hold in Algorithm 2.
@@
Under Assumption 7.1 and 7.3, the iterates $\{(x_{k},\pi_{k},w_{k})\}$ generated by Algorithm 2 satisfy
$\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}\leq$ $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
@@
By the update rule of $w$, $w_{k}=w_{k-1}-\eta\sigma_{k}$, we decompose the first term in (32):
$\left\|w_{k}-w_{k-1}^{*}\right\|_{2}^{2}=\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}-2\eta\left<\sigma_{k},w_{k-1}-w_{k-1}^{*}\right>.$ (33)
@@
$\eta C_{\sigma\pi}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}\leq$ $\frac{\eta^{2}}{2\rho_{1}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}+\frac{C_{\sigma\pi}^{2}\rho_{1}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}$
$=$ $\frac{\eta^{2}C_{\sigma\pi}^{2}}{2\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}+\frac{\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$
@@
$\eta\left<\sigma_{k},w_{k-1}-w_{k-1}^{*}\right>\geq\frac{\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}-\frac{\eta^{2}C_{\sigma\pi}^{2}}{2\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}.$ (35)
@@
$\left\|w_{k}-w_{k-1}^{*}\right\|_{2}^{2}\leq\left(1-\eta\left(1-\gamma\right)^{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\frac{\eta^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}.$ (36)
@@
$\leq$ $\ \rho_{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{1}{\rho_{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}$
$=$ $\ \frac{\eta\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{2}{\eta\left(1-\gamma\right)^{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}$
$+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2},$ (38)
where we take $\,\,\rho_{2}=\frac{\eta\left(1-\gamma\right)^{2}}{2}$. Assembling (36), (37) and (38), we can estimate $\left\|w_{k}-w_{k}^{*}\right\|_{2}^{2}$ in (32),
@@
$\leq$ $\left(1-\eta\left(1-\gamma\right)^{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\frac{\eta^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+\left\|w_{k-1}^{*}-w_{k}^{*}\right\|_{2}^{2}+\frac{\eta\left(1-\gamma\right)^{2}}{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}$
$+\frac{2}{\eta\left(1-\gamma\right)^{2}}\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}+\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}\,\,$
$=$ $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\frac{\eta^{2}C_{\sigma\pi}^{2}}{\left(1-\gamma\right)^{2}}\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\left\|w_{k}^{*}-w_{k-1}^{*}\right\|_{2}^{2}+2\eta^{2}\left\|\sigma_{k}\right\|_{2}^{2}$
$\leq$ $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}\right)\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2}+\eta^{2}C_{\sigma\pi}^{2}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\left\|\pi_{k}-\pi_{k}^{*}\right\|_{\infty}^{2}$
$+2L_{w}^{2}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\left\|x_{k}-x_{k-1}\right\|_{2}^{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\left\|w_{k-1}-w_{k-1}^{*}\right\|_{2}^{2},$